OpenStudy (anonymous):

Evaluate: lim_{n rightarrow infty}sum_{k=1}^{n}(-1+9k^{2}div n ^{2})*3div n I just need help with the steps. (PLEEEZZEE)

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}\sum_{k=1}^{n}(-1+9k^{2}\div n ^{2})*3\div n\]

OpenStudy (anonymous):

and how do I evaluate that?

OpenStudy (anonymous):

I think it's best to first see if it is convergent :-D

OpenStudy (jamesj):

First find an expression for \( \sum_1^n k^2 \). Hence you can eliminate the sum and k from the expression altogether and just have an expression in n.

OpenStudy (anonymous):

I think there is an expression for that, but I forgot what it was :-P let's google

OpenStudy (jamesj):

I'll tell you right now that \[ \sum_1^n k^2 = \frac{1}{6}n(n+1)(2n+1) \] Using that, eliminate k from your expression.

OpenStudy (anonymous):

\[\frac{n(n+1(2n+1)}{6}\]

OpenStudy (anonymous):

aww beat me to it :) alright how do we fit that ugly fraction into that

OpenStudy (anonymous):

I think we should break that sum apart

OpenStudy (jamesj):

\[ \sum_{k=1}^n \frac{-1+9k^2 }{n^2} \frac{3}{n} = \frac{3}{n^3} \left( -\sum_{k=1}^n 1 + \sum_{k=1}^n 9k^2 \right) \] Now you kwenisha, take it from here.

OpenStudy (anonymous):

so no we have: \[ 3divn(\sum_{k=1}^{n}-1+9divn \sum_{k=1}^{n}k ^{2}) => 3divn(-n+3divn(n+1)(2n+1)\div2\], this then give us: \[3divn(-n+3(n+1)(2n+1))\div2n\] and on and on and on... think I got the general idea :)

OpenStudy (anonymous):

So *now* we have... not *no* :)

OpenStudy (jamesj):

Be careful. You've already at least one mistake I can see.

OpenStudy (anonymous):

oops...

OpenStudy (anonymous):

Quick question: is this the general formula/equation when solving a problem like this n(n+1(2n+1)/6?